# MEGA: Memory-Efficient 4D Gaussian Splatting for Dynamic Scenes

### v1

@@ 4.3 Ablation Study
To validate the effectiveness of various components within our proposed method, we conduct ablation experiments on selected scenes from two datasets. We analyze the impact of these components on scenes from the Technicolor dataset (Birthday, Fabien) and the Neu3DV dataset (Flame Steak, Sear Steak). Detailed results are presented in Table 3.
Compact DC-AC Color Representation. Building on the original 4DGS, we substitute the 4D SH coefficients with a grid-based neural field representation (Lee et al., 2024), and our proposed DAC representation, respectively. While the grid-based approach, referred to as “w/ grid,” achieves a reduction of approximately 10$\times$ in parameters, it leads to a significant performance degradation compared to 4DGS. This performance loss may be attributed to the grid’s inability to retain sufficient detail, thereby discarding critical information. To address this issue, we use a DC component to preserve essential color information inherently present in the scene, and an AC predictor to encode the temporal-viewpoint variations in color. This method allows us to achieve a comparable reduction in storage as the grid-based approach while maintaining high-quality rendering consistent with 4DGS.
Entropy-constrained Gaussian Deformation. This part of our ablation study evaluates the impact of Gaussian deformation and opacity-based entropy loss $\mathcal{L}_{opa}$. Starting from the configuration “w/ DAC”, we observe that implementing a deformation predictor alone (referred to as “w/ DAC+Deformation”) leads to an increased number of Gaussians. Conversely, employing $\mathcal{L}_{opa}$ without the deformation predictor (referred to as “w/ DAC+$\mathcal{L}_{opa}$”) limits the action range of each Gaussian, inhibiting their efficacy. However, when combining our deformation predictor with $\mathcal{L}_{opa}$, this strategy significantly reduces the number of Gaussians needed while maintaining rendering quality comparable to that of 4DGS.
Figure 5: Subjective comparison of various methods on Theater scene (Top) and Painter scene (Bottom) from the Technicolor Dataset.
@@ Appendix B Network Structure
AC Color Predictor. Fig. 8 (a) shows the details of the AC color predictor. After generating the AC color component $\bm{c}_{ac}^{t,v}$, we combine the DC component $\bm{c}_{dc}$ to produce the final color $\bm{c}_{t,v}$.
@@
DyNeRF 29.20 - 0.0240 0.0668 - - 32.76 - 0.0175 0.2417 - -
HyperReel 29.99 0.0390 - 0.0531 - - 34.70 0.0525 - 0.1864 - -
Deformable 3DGS 30.68 0.0440 0.0237 0.0775 52.83 90.61MB 33.33 0.0673 0.0273 0.1851 95.52 42.81MB
E-D3DGS 31.88 0.0328 0.0172 0.0506 62.41 66.50MB 34.69 0.0612 0.0236 0.1689 124.71 20.02MB
STG 31.65 0.0293 0.0156 0.0413 128.43 51.81MB 35.61 0.0468 0.0177 0.1140 138.03 40.23MB
4DGS 31.00 0.0383 0.0211 0.0629 39.61 7986.31MB 33.57 0.0582 0.0226 0.1555 87.54 3334.57MB
@@
DyNeRF 35.95 - 0.0140 0.1464 - - 29.53 - 0.0305 0.1881 - -
HyperReel 35.91 0.0385 - 0.1173 - - 33.32 0.0525 - 0.1154 - -
Deformable 3DGS 34.71 0.0497 0.0211 0.1302 84.37 51.56MB 29.65 0.0768 0.0382 0.1795 80.40 54.75MB
E-D3DGS 35.97 0.0360 0.0149 0.0903 94.91 38.00MB 31.04 0.0643 0.0307 0.1493 56.88 77.61MB
STG 35.73 0.0369 0.0148 0.0963 157.01 54.84MB 31.16 0.0595 0.0286 0.1332 137.48 48.52MB
4DGS 35.73 0.0423 0.0176 0.1125 54.73 5667.79MB 31.29 0.0696 0.0341 0.1653 54.05 5770.69MB
@@
DyNeRF 31.58 - 0.0190 0.0670 - - 31.80 - 0.0210 0.1400 0.02 30.00MB
HyperReel 29.74 0.0525 - 0.0723 - - 32.70 0.0470 - 0.1090 4.00 60.00MB
Deformable 3DGS 26.39 0.1104 0.0663 0.2040 67.32 67.08MB 30.95 0.0696 0.0353 0.1553 76.09 61.36MB
E-D3DGS 30.87 0.0525 0.0289 0.0976 56.81 78.23MB 32.89 0.0494 0.0231 0.1114 79.14 56.07MB
STG 32.61 0.0296 0.0169 0.0380 147.70 61.34MB 33.35 0.0404 0.0187 0.0846 141.73 51.35MB
4DGS 28.79 0.0590 0.0362 0.0985 40.36 7775.97MB 32.07 0.0535 0.0263 0.1189 55.26 6107.07MB
@@
HexPlane2,3 - - - - - - 32.04 - 0.0150 0.0820 - -
NeRFPlayer3 31.53 0.0245 - 0.085 - - 30.56 0.0355 - 0.1130 - -
HyperReel 28.37 0.0540 - 0.1270 - - 32.30 0.0295 - 0.0890 - -
K-Planes 29.99 - 0.0170 - - - 31.82 - 0.0170 - - -
MixVoxels-L 29.63 - 0.0162 0.099 - - 32.40 - 0.0157 0.088 - -
MixVoxels-X 30.39 - 0.0160 0.062 - - 32.63 - 0.0146 0.057 - -
Dynamic 3DGS 26.49 0.0263 0.0129 0.087 - - 30.72 0.0295 0.0161 0.090 - -
Deformable 3DGS 27.88 0.0470 0.0284 0.0855 26.89 33.84MB 33.06 0.0267 0.0142 0.0519 31.06 33.21MB
E-D3DGS 29.56 0.0319 0.0193 0.0300 51.94 57.97MB 32.71 0.0219 0.0123 0.0255 74.11 36.82MB
STG3 28.55 0.0418 0.0253 0.0692 221.76 214.52MB 33.18 0.0215 0.0113 0.0367 290.03 151.52MB
4DGS 27.98 0.0435 0.0265 0.0847 78.79 3704.58MB 32.73 0.0245 0.0133 0.0489 111.77 2474.94MB
@@
Neural Volume1 - - - - - - 22.80 - 0.0620 0.2950 - -
LLFF1 - - - - - - 23.24 - 0.0200 0.2350 - -
DyNeRF1 - - - - - - 29.58 - 0.0200 0.0830 0.015 28.00MB
HexPlane2,3 32.55 - 0.0130 $0$ 29.47 - 0.0180 0.0780 - -
NeRFPlayer3 29.35 0.0460 - 0.1440 - - 31.65 $0$ 0.098 - -
HyperReel 32.92 0.0275 - 0.084 - - 28.26 0.0590 - 0.136 - -
K-Planes 31.82 - 0.0170 - - - 30.44 - 0.0235 - - -
MixVoxels-L 32.40 - 0.0157 0.088 - - 29.81 - 0.0255 0.116 - -
MixVoxels-X 32.63 - 0.0146 0.057 - - 30.60 - 0.0233 0.078 - -
Dynamic 3DGS 30.72 0.0295 0.0161 $0$ 26.92 0.0512 0.0302 0.1220 - -
Deformable 3DGS 31.43 0.0333 0.0204 0.0551 28.43 33.14MB 28.70 0.0432 0.0255 0.0804 28.72 34.17MB
E-D3DGS 33.02 0.0213 0.0116 0.0258 74.33 36.63MB 29.79 0.0363 0.0216 0.0535 61.03 45.08MB
STG3 33.55 0.0207 0.0106 0.0367 299.98 135.28MB 29.48 0.0375 0.0224 0.0630 215.69 268.39MB
4DGS 33.23 0.0226 0.0119 0.0470 109.11 2555.56MB 28.86 0.0425 0.0257 0.0832 64.31 4695.46MB
@@
HexPlane2,3 32.08 - 0.0110 0.0660 - - 32.39 - 0.0110 $0$
NeRFPlayer3 31.93 0.0250 - 0.0880 - - 29.13 0.0460 - 0.138 - -
HyperReel 32.20 0.0255 - 0.078 - - 32.57 0.0240 - 0.077 - -
K-Planes 32.38 - 0.0150 - - - 32.52 - 0.0130 - - -
MixVoxels-L 31.83 - 0.0144 0.088 - - 32.10 - 0.0122 0.080 - -
MixVoxels-X 32.10 - 0.0137 0.051 - - 32.33 - 0.0121 0.053 - -
Dynamic 3DGS 33.24 0.0233 0.0113 0.0790 - - 33.68 0.0224 0.0105 0.079 - -
Deformable 3DGS 31.83 0.0248 0.0137 0.0418 30.91 30.72MB 33.01 0.0237 0.0125 0.0416 31.73 30.74MB
E-D3DGS 30.23 0.0241 0.0149 0.0243 76.92 32.244MB 31.91 0.0200 0.0110 0.0233 79.89 32.426MB
STG3 33.59 0.0178 0.0088 0.0290 305.22 141.25MB 33.89 0.0174 0.0085 0.0295 308.15 141.16MB
4DGS 33.19 0.0204 0.0106 0.0389 91.52 3173.37MB 33.44 0.0204 0.0105 0.0411 124.66 2164.07MB
@@
Neural Volume1 22.80 - 0.0620 0.2950 - -
LLFF1 23.24 - 0.0200 0.2350 - -
DyNeRF1 29.58 - 0.0200 0.0830 0.015 28.00MB
HexPlane2,3 31.71 - 0.0140 0.0750 0.56 200.00MB
StreamRF4 28.26 - - - 10.90 5310.00MB
NeRFPlayer3 30.69 0.0340 - 0.1110 0.05 5130.00MB
HyperReel 31.10 0.0360 - 0.0960 2.00 360.00MB
K-Planes 31.63 - 0.0180 - 0.30 311.00MB
MixVoxels-L 31.34 - 0.0170 0.0960 37.70 500.00MB
MixVoxels-X 31.73 - 0.0150 0.0640 4.60 500.00MB
Dynamic 3DGS 30.46 0.0350 0.0190 0.0990 460.00 2772.00MB
C-D3DGS4 30.46 - - 0.1500 118.00 338.00MB
Deformable 3DGS 30.98 0.0331 0.0191 0.0594 29.62 32.64MB
E-D3DGS 31.20 0.0259 0.0151 0.0304 69.70 40.20MB
STG3 32.04 0.0261 0.0145 0.0440 273.47 175.35MB
4DGS 31.57 0.0290 0.0164 0.0573 96.69 3128.00MB
@@
Figure 6: Subjective comparison of various methods on Birthday scene (Top) and Trains scene (Bottom) from the Technicolor Dataset.

### v2

@@ 4.3 Ablation Study
To validate the effectiveness of various components within our proposed method, we conduct ablation experiments on on scenes from the Technicolor dataset (Birthday, Fabien) and the Neu3DV dataset (Flame Steak, Sear Steak). Detailed results are presented in Table 3.
Compact DC-AC Color Representation. Building on the original 4DGS, we substitute the 4D SH coefficients with a grid-based neural field representation [20], and our proposed DAC representation, respectively. While the grid-based approach, referred to as “w/ grid," achieves a reduction of approximately 10$\times$ in parameters, it leads to a significant performance degradation compared to 4DGS. This performance loss may be attributed to the grid’s inability to retain sufficient detail, thereby discarding critical information. To address this issue, we use a DC component to preserve essential color information inherently present in the scene, and an AC predictor to encode the temporal-viewpoint variations in color. This method allows us to achieve a comparable reduction in storage as the grid-based approach while maintaining high-quality rendering consistent with 4DGS.
Entropy-constrained Gaussian Deformation. This part of our ablation study evaluates the impact of Gaussian deformation and opacity-based entropy loss $\mathcal{L}_{opa}$. Starting from the configuration “w/ DAC”, we observe that implementing a deformation predictor alone (referred to as “w/ DAC+Deformation”) leads to an increased number of Gaussians. Conversely, employing $\mathcal{L}_{opa}$ without the deformation predictor (referred to as “w/ DAC+$\mathcal{L}_{opa}$") limits the action range of each Gaussian, inhibiting their efficacy. However, when combining our deformation predictor with $\mathcal{L}_{opa}$, this strategy significantly reduces the number of Gaussians needed while maintaining rendering quality comparable to that of 4DGS.
@@ 5 Conclusion
Acknowledgments. This work was supported by the Hong Kong Research Grants Council under the Areas of Excellence scheme grant AoE/E-601/22-R and NSFC/RGC Collaborative Research Scheme grant CRS_HKUST603/22.
Supplementary Material
@@ Appendix B Network Structure
AC Color Predictor. Fig. 8 (a) shows the details of the AC color predictor. After generating the AC color component $\boldsymbol{c}_{ac}^{t,v}$, we combine the DC component $\boldsymbol{c}_{dc}$ to produce the final color $\boldsymbol{c}_{t,v}$.
@@
Figure 7: Subjective comparison of various methods on Birthday scene (Top), Trains scene (Medium) and Painter scene (Bottom) from the Technicolor Dataset.
@@
DyNeRF [22] 29.20 - 0.0240 0.0668 - - 32.76 - 0.0175 0.2417 - -
HyperReel [1] 29.99 0.0390 - 0.0531 - - 34.70 0.0525 - 0.1864 - -
Deformable 3DGS [43] 30.68 0.0440 0.0237 0.0775 52.83 90.61MB 33.33 0.0673 0.0273 0.1851 95.52 42.81MB
E-D3DGS [2] 31.88 0.0328 0.0172 0.0506 62.41 66.50MB 34.69 0.0612 0.0236 0.1689 124.71 20.02MB
STG [23] 31.65 0.0293 0.0156 0.0413 128.43 51.81MB 35.61 0.0468 0.0177 0.1140 138.03 40.23MB
4DGS [45] 31.00 0.0383 0.0211 0.0629 39.61 7986.31MB 33.57 0.0582 0.0226 0.1555 87.54 3334.57MB
@@
DyNeRF [22] 35.95 - 0.0140 0.1464 - - 29.53 - 0.0305 0.1881 - -
HyperReel [1] 35.91 0.0385 - 0.1173 - - 33.32 0.0525 - 0.1154 - -
Deformable 3DGS [43] 34.71 0.0497 0.0211 0.1302 84.37 51.56MB 29.65 0.0768 0.0382 0.1795 80.40 54.75MB
E-D3DGS [2] 35.97 0.0360 0.0149 0.0903 94.91 38.00MB 31.04 0.0643 0.0307 0.1493 56.88 77.61MB
STG [23] 35.73 0.0369 0.0148 0.0963 157.01 54.84MB 31.16 0.0595 0.0286 0.1332 137.48 48.52MB
4DGS [45] 35.73 0.0423 0.0176 0.1125 54.73 5667.79MB 31.29 0.0696 0.0341 0.1653 54.05 5770.69MB
@@
DyNeRF [22] 31.58 - 0.0190 0.0670 - - 31.80 - 0.0210 0.1400 0.02 30.00MB
HyperReel [1] 29.74 0.0525 - 0.0723 - - 32.70 0.0470 - 0.1090 4.00 60.00MB
Deformable 3DGS [43] 26.39 0.1104 0.0663 0.2040 67.32 67.08MB 30.95 0.0696 0.0353 0.1553 76.09 61.36MB
E-D3DGS [2] 30.87 0.0525 0.0289 0.0976 56.81 78.23MB 32.89 0.0494 0.0231 0.1114 79.14 56.07MB
STG [23] 32.61 0.0296 0.0169 0.0380 147.70 61.34MB 33.35 0.0404 0.0187 0.0846 141.73 51.35MB
4DGS [45] 28.79 0.0590 0.0362 0.0985 40.36 7775.97MB 32.07 0.0535 0.0263 0.1189 55.26 6107.07MB
@@
HexPlane2,3 [3] - - - - - - 32.04 - 0.0150 0.0820 - -
NeRFPlayer3 [38] 31.53 0.0245 - 0.085 - - 30.56 0.0355 - 0.1130 - -
HyperReel [1] 28.37 0.0540 - 0.1270 - - 32.30 0.0295 - 0.0890 - -
K-Planes [13] 29.99 - 0.0170 - - - 31.82 - 0.0170 - - -
MixVoxels-L [40] 29.63 - 0.0162 0.099 - - 32.40 - 0.0157 0.088 - -
MixVoxels-X [40] 30.39 - 0.0160 0.062 - - 32.63 - 0.0146 0.057 - -
Dynamic 3DGS [28] 26.49 0.0263 0.0129 0.087 - - 30.72 0.0295 0.0161 0.090 - -
Deformable 3DGS [43] 27.88 0.0470 0.0284 0.0855 26.89 33.84MB 33.06 0.0267 0.0142 0.0519 31.06 33.21MB
E-D3DGS [2] 29.56 0.0319 0.0193 0.0300 51.94 57.97MB 32.71 0.0219 0.0123 0.0255 74.11 36.82MB
STG3 [23] 28.55 0.0418 0.0253 0.0692 221.76 214.52MB 33.18 0.0215 0.0113 0.0367 290.03 151.52MB
4DGS [45] 27.98 0.0435 0.0265 0.0847 78.79 3704.58MB 32.73 0.0245 0.0133 0.0489 111.77 2474.94MB
@@
Neural Volume1 [25] - - - - - - 22.80 - 0.0620 0.2950 - -
LLFF1 [29] - - - - - - 23.24 - 0.0200 0.2350 - -
DyNeRF1 [22] - - - - - - 29.58 - 0.0200 0.0830 0.015 28.00MB
HexPlane2,3 [3] 32.55 - 0.0130 $0$ 29.47 - 0.0180 0.0780 - -
NeRFPlayer3 [38] 29.35 0.0460 - 0.1440 - - 31.65 $0$ 0.098 - -
HyperReel [1] 32.92 0.0275 - 0.084 - - 28.26 0.0590 - 0.136 - -
K-Planes [13] 31.82 - 0.0170 - - - 30.44 - 0.0235 - - -
MixVoxels-L [40] 32.40 - 0.0157 0.088 - - 29.81 - 0.0255 0.116 - -
MixVoxels-X [40] 32.63 - 0.0146 0.057 - - 30.60 - 0.0233 0.078 - -
Dynamic 3DGS [28] 30.72 0.0295 0.0161 $0$ 26.92 0.0512 0.0302 0.1220 - -
Deformable 3DGS [43] 31.43 0.0333 0.0204 0.0551 28.43 33.14MB 28.70 0.0432 0.0255 0.0804 28.72 34.17MB
E-D3DGS [2] 33.02 0.0213 0.0116 0.0258 74.33 36.63MB 29.79 0.0363 0.0216 0.0535 61.03 45.08MB
STG3 [23] 33.55 0.0207 0.0106 0.0367 299.98 135.28MB 29.48 0.0375 0.0224 0.0630 215.69 268.39MB
4DGS [45] 33.23 0.0226 0.0119 0.0470 109.11 2555.56MB 28.86 0.0425 0.0257 0.0832 64.31 4695.46MB
@@
HexPlane2,3 [3] 32.08 - 0.0110 0.0660 - - 32.39 - 0.0110 $0$
NeRFPlayer3 [38] 31.93 0.0250 - 0.0880 - - 29.13 0.0460 - 0.138 - -
HyperReel [1] 32.20 0.0255 - 0.078 - - 32.57 0.0240 - 0.077 - -
K-Planes [13] 32.38 - 0.0150 - - - 32.52 - 0.0130 - - -
MixVoxels-L [40] 31.83 - 0.0144 0.088 - - 32.10 - 0.0122 0.080 - -
MixVoxels-X [40] 32.10 - 0.0137 0.051 - - 32.33 - 0.0121 0.053 - -
Dynamic 3DGS [28] 33.24 0.0233 0.0113 0.0790 - - 33.68 0.0224 0.0105 0.079 - -
Deformable 3DGS [43] 31.83 0.0248 0.0137 0.0418 30.91 30.72MB 33.01 0.0237 0.0125 0.0416 31.73 30.74MB
E-D3DGS [2] 30.23 0.0241 0.0149 0.0243 76.92 32.244MB 31.91 0.0200 0.0110 0.0233 79.89 32.426MB
STG3 [23] 33.59 0.0178 0.0088 0.0290 305.22 141.25MB 33.89 0.0174 0.0085 0.0295 308.15 141.16MB
4DGS [45] 33.19 0.0204 0.0106 0.0389 91.52 3173.37MB 33.44 0.0204 0.0105 0.0411 124.66 2164.07MB
@@
Neural Volume1 [25] 22.80 - 0.0620 0.2950 - -
LLFF1 [29] 23.24 - 0.0200 0.2350 - -
DyNeRF1 [22] 29.58 - 0.0200 0.0830 0.015 28.00MB
HexPlane2,3 [3] 31.71 - 0.0140 0.0750 0.56 200.00MB
StreamRF4 [21] 28.26 - - - 10.90 5310.00MB
NeRFPlayer3 [38] 30.69 0.0340 - 0.1110 0.05 5130.00MB
HyperReel [1] 31.10 0.0360 - 0.0960 2.00 360.00MB
K-Planes [13] 31.63 - 0.0180 - 0.30 311.00MB
MixVoxels-L [40] 31.34 - 0.0170 0.0960 37.70 500.00MB
MixVoxels-X [40] 31.73 - 0.0150 0.0640 4.60 500.00MB
Dynamic 3DGS [28] 30.46 0.0350 0.0190 0.0990 460.00 2772.00MB
C-D3DGS4 [16] 30.46 - - 0.1500 118.00 338.00MB
Deformable 3DGS [43] 30.98 0.0331 0.0191 0.0594 29.62 32.64MB
E-D3DGS [2] 31.20 0.0259 0.0151 0.0304 69.70 40.20MB
STG3 [23] 32.04 0.0261 0.0145 0.0440 273.47 175.35MB
4DGS [45] 31.57 0.0290 0.0164 0.0573 96.69 3128.00MB
@@
Appendix C Ablation Study
MLP. As shown in Table 6, MLPs of various sizes exhibit similar results, because $\mathcal{F}_{\boldsymbol{\phi}}$ and $\mathcal{F}_{\boldsymbol{\theta}}$ only provide temporal and viewpoint varying information, which can be effectively captured by lightweight MLPs.
Trade-off coefficients. Table 6 shows the results of various trade-off coefficients $\lambda$ and $\kappa$. Our default trade-off coefficients are chosen empirically.
Opacity loss. As shown in Table 6, simply applying $\mathcal{L}_{opa}$ to existing baselines does not improve performance. For 4DGS and STG, the reason may arise from their explicit modeling of motion as fixed low-order polynomials, e.g., linear or quadratic. Thus, enforcing sparsity in opacity may conflict with the motion priors, resulting in either over-pruning or insufficient flexibility to model more complex nonlinear temporal dynamics. For E-D3DGS, while it supports more flexible motion via multi-granularity embeddings, it induces locally similar deformation by regularizing nearby per-Gaussian embeddings. Therefore, adding $\mathcal{L}_{opa}$ may disrupt this smooth deformation by encouraging abrupt temporal activation, degrading local coherence. In contrast, our method jointly learns deformation and opacity in a unified way, enabling both localized adaptation and temporal sparsity without conflict.
View direction. Since the same time frames may reveal different visible content under different viewpoints, the view direction input is critical for disambiguating view-dependent geometry and motion, especially in sparse or occluded camera settings. As shown in Table 7, removing the view input significantly degrades rendering quality and requires more Gaussians, indicating less efficient and less accurate scene modeling.
Table 6: Ablation study on the Fabien scene. $\mathcal{F}_{\boldsymbol{\phi}}$ denotes the color MLP network, $\mathcal{F}_{\boldsymbol{\theta}}$ denotes the deformation MLP network. The first row denotes our final solution with $\lambda=0.2$ and $\kappa=5e^{-4}$.
[…]
Ours 34.89 0.0597 0.31M 6.43M
Large $\mathcal{F}_{\boldsymbol{\phi}}$ 33.59 0.0653 0.35M 7.30M
Large $\mathcal{F}_{\boldsymbol{\theta}}$ 34.71 0.0604 0.33M 8.24M
Large $\mathcal{F}_{\boldsymbol{\phi}}$+$\mathcal{F}_{\boldsymbol{\theta}}$ 34.27 0.0627 0.30M 7.85M
$\lambda$=0.1, $\kappa$=$5e^{-4}$ 34.09 0.0643 0.27M 5.74M
$\lambda$=0.3, $\kappa$=$5e^{-4}$ 33.99 0.0627 0.47M 9.77M
$\lambda$=0.2, $\kappa$=$1e^{-4}$ 33.99 0.0639 0.48M 10.00M
$\lambda$=0.2, $\kappa$=$1e^{-3}$ 34.01 0.0633 0.31M 6.41M
4DGS [45] 33.57 0.0582 5.43M 874.14M
4DGS+$\mathcal{L}_{opa}$ 23.23 0.1037 8.41M 1353.13M
STG [23] 35.61 0.0468 0.30M 10.54M
STG+$\mathcal{L}_{opa}$ 33.78 0.0610 0.28M 26.03M
E-D3DGS [2] 34.69 0.0612 0.06M 5.25M
E-D3DGS+$\mathcal{L}_{opa}$ 34.52 0.0623 0.11M 10.21M
Table 7: Effect of view direction on the Birthday scene.
[…]
Ours 32.02 0.0309 0.91M 18.48M
w/o view 27.35 0.0697 1.54M 31.44M
Appendix D Limitations
First, MEGA lacks robustness to real-world noise such as motion blur and color inconsistency. Integrating techniques from Robust GS [8] is a promising future direction. Second, while MEGA achieves significant compression, the increased flexibility of 4D Gaussians can lead to artifacts in cases of ultra-fast motion. A potential solution is to decompose the scene into static and dynamic regions, and then use 3D Gaussians for static regions and 4D Gaussians for dynamic regions. Finally, MEGA struggles with novel view extrapolation beyond the bound of training views, which can introduce artifacts. Incorporating strong scene priors could help improve generalization.